the perimeter of a rectangle is 44inches.find the dimensions if the width is two inches longer than three times the length.hen find the area of the rectangle.
Answer by checkley71(8403)

W=3L+2
2L+2(3L+2)=44
2L+6L+4=44
8L=44-4
8L=40
L=40/8
L=5 ANSWER FOR THE LENGTH.
W=3*5+2
W=15+2
W=17 ANSWER FOR THE WIDTH.
PROOF:
2*5+2*17=44
10+34=44
44=44
AREA=5*17
AREA=85 ANSWER.